# On the Presence of Angular-Velocity Offsets in Disk Galaxies

### v1

@@ Abstract
The well-known discrepancy in galactic rotation curves refers to the mismatch between observed rotational velocities and the velocities predicted by baryonic matter. In this study we investigate a potential pattern in the discrepancy, which may point to an underlying pattern in dark-halo distributions. By looking at rotational-velocity curves from an alternate perspective, the angular-velocity curves, it appears that the observed angular velocities and their corresponding baryonic predictions differ by a constant shift. That is, the discrepancy may be reduced to a constant angular-velocity term, independent of the radius. We test the generality of the suggested property by analysing 143 high-quality rotation curves. The property appears significant as it performs equally well (or better) than well-established dark-halo models. Compared to a Burkert profile, it is preferred in 60% of the cases, while relative to a Navarro–Frenk–White profile (NFW), it is superior in 73% of the cases. Next, by including the new phenomenological property within the dynamical equations, we find an explicit expression for the dark-halo profile. The new single-parameter profile is characterised by a remarkable property: it is intrinsically related to the baryonic distribution. Thus, information regarding the cuspy or cored nature of a particular dark halo, according to this profile, is encoded (and explicitly determined) by the respective baryonic behaviour.
@@
pubyear: 2025pagerange: On the Presence of Angular-Velocity Offsets in Disk GalaxiesB
@@ 1 Introduction
One of the most useful tools to analyse the dynamics of disk galaxies and derive their mass distributions is a galactic rotation curve (Bosma, 1981; Sofue & Rubin, 2001). A rotation curve (RC) presents the circular velocities in a galaxy as a function of the distance from the galaxy’s center. Historically, the discrepancy revealed in RCs is known as the first robust evidence for dark matter (Babcock, 1939; Salpeter, 1978; Rubin et al., 1978, 1980; Carignan, 1985; Sancisi, 1987). Illustrations of the discrepancy (i.e. the mismatch between baryonic predictions and observations) are shown in the upper panels of Fig. 1.
@@
Although the baryonic distributions are less dominant than their accompanying dark halos (Kuhlen et al., 2012; Reddick et al., 2013), their impact on the overall dynamics seems to be important. A notable result of such a kind was found by Sancisi (2004), who showed that features in observed baryonic distributions closely correlate with features in rotation curves. Since the velocities in rotation curves reflect the total gravitational field (i.e., of the overall mass distribution), such a correlation highlights the important role of baryons in describing the total dynamics.
@@
The objective of the current paper is to demonstrate, establish, and explore the implications of the constant-offsets property. A significant outcome of the property is an explicit relation that connects dark and baryonic distributions in disk galaxies. This can be achieved by phenomenologically integrating the offsets into the dynamics. The resulting relation can be interpreted as a new data-driven dark-halo profile.
@@
The next sections are organized as follows: Section 2 presents the new concept of angular-velocity curves and explores the universality of constant offsets. Section 3 calculates the baryonic masses of 143 disk galaxies using their observed rotation curves, and without additional assumptions. Section 4 derives a relation that connects baryonic distributions with their host dark halos, or equivalently, introduces a new dark-halo profile. In section 5, we discuss potential interpretations of the observed property, conclude, and propose future directions.
@@ 2 Angular-Velocity Curves
Let us now dive into the analysis and explore the universality of the offsets. For this purpose we use the publicly-available SPARC database (Lelli et al., 2016) together with the catalog of SPARC dark-halo models (Li et al., 2020). SPARC is a database that includes 175 nearby galaxies with new surface photometry at 3.6 $\mu$m and high-quality rotation curves. The publicly available catalog of Li et al. (2020) consists of 175 rotation-curve fits (for the entire SPARC sample) using seven different dark-halo profiles.
@@
Our aim is to compare the new phenomenological relation (Eq. 1) with established approaches. Since models that include standard dark-halo profiles already produce reliable fits to the data, they may act as a baseline. If the new model (baryons + offset) could produce comparable results to standard models, it may be considered a viable approach. This would support the universality of angular-velocity offsets in disk galaxies.
@@
Figure 3: A visual representation of the sample’s best-fit model parameters. Left panel: A scatter plot of the two parameters. There is a weak but not significant anti-correlation between them ($\rm\rho_{pearson}=-0.1$, $p_{\rm value}=0.26$). Middle panel: Histogram of M/L values. The median value of log(M/L) is 0.06 and its standard error is 0.43. There are four outliers in the data with $\rm M/L<0.1$, but within uncertainties, they are also consistent with higher values. Right Panel: Histogram of $\rm\omega_{0}$ values. The median value of $\rm log(\omega_{0})$ is -15.97 and its standard error is 0.51. Eight points with $\rm\omega_{0}<0$ are not shown, due to the logarithmic nature of the plot, but within uncertainties, they are also consistent with positive values.
Dark-halo models usually include three free parameters: $\rm M/L$ for the baryonic component and two more for the dark halo. In the current analysis we focus on two dark halo profiles: the cosmology-driven NFW (Navarro et al., 1997) and the data-driven Burkert (Burkert, 1995). In order to quantify the goodness of fit of the dark-halo models we rely on the catalog of Li et al. (2020). For each galactic rotation curve, the catalog specifies the best-fit parameters and the corresponding $\chi^{2}$.
Since we prefer to use exactly the same data-sets when testing the competing models (i.e., NFW, Burkert and the current one), we convert the angular velocities of Eq.  1 into actual velocities, as follows:
@@
where $\rm M/L$ is the stellar-disk mass-to-light ratio, $v_{\rm disk}(r)$ is the normalized stellar-disk contribution and $v_{\rm gas}(r)$ is the gaseous-disk contribution. The normalized stellar-disk and the gaseous-disk contributions are included in the SPARC database as data files. They were numerically derived from the corresponding light and neutral hydrogen distributions. More information on the process can be found in Sanders & McGaugh (2002). For simplicity, only galaxies with negligible bulge contributions were selected in this study, which left us with 143 galaxies.
The next step is to apply priors to the parameters. For the $M/L$ parameter we impose a prior of the form $\rm M/L$ > 0 to ensure that it maintains its physical meaning. For $\rm\omega_{0}$ we do not apply a prior since no earlier knowledge is available. We then run the simulation to produce the fittings, estimate the parameters’ values, and extract the corresponding $\rm\chi^{2}$ of each fit. An example of an MCMC output (converted to angular velocities for convenience) is given in Fig. 2.
@@
Another aspect would be the typical values of the parameters. Fig. 3 displays the distributions of the parameters. Table 1 provides the values, their errors, and additional statistical information. The median value of $\rm\omega_{0}$ is $\sim 10^{-16}$ [rad/s], with a range of almost two orders of magnitude. The median value of $\rm M/L$ is $\sim\rm 1$, which is a bit larger than expectations (McGaugh & Schombert, 2014). There is a weak but not significant anti-correlation between the parameters ($\rm\rho_{pearson}=-0.1$, $p_{\rm value}=0.26$).
Table 1: Best-fit parameters and fitting information.
Galaxy $\rm M/L$ [$\rm M_{\odot}/L_{\odot}$] $\rm\omega_{0}$ $\rm[10^{-16}\,rad/s]$ $n$ $\rm\chi^{2}$ $\rm p_{value}$
@@
With both the MCMC and the catalog results in hand, the relative performance of the different models could be quantified. That is, the performance of the phenomenological model with respect to the standard dark halo models. For this purpose we employ the Bayesian Information Criterion (BIC). BIC is a criterion for model selection that accounts for the trade-off between a model’s fit and its complexity. It is suitable for comparing non-nested models when the data are assumed to be independent and normally distributed, with the noise being well characterized by measurement errors. The BIC score of a given model (with respect to a particular dataset) can be written as:
@@
On the right side of Fig. 4, we plot a similar histogram, with the comparison being relative to Burkert’s profile. In this case, the $\rm\Delta BIC$ test suggests that the models are equally credible, with slightly better results for the phenomenological model (median value of 0.77 $\pm$ 0.36). Flawed results were excluded as well (14 galaxies where Burkert fails, 20 where the phenomenological model fails, and 16 where both models fail).
While the number of excluded galaxies appears to be high, the essential aspect here is that the numbers are comparable. In some cases the standard dark halo models fail, whereas in others it is the phenomenological model that has difficulty fitting the data. The challenges in fitting the data depend on the accuracy of the data itself but also on the modeling of the baryonic component. These difficulties may include the extraction of surface photometry (Lelli et al., 2016), the bulge-disk decomposition (Simard et al., 2011) and the irregular morphologies of low surface-brightness galaxies.
Another way to look at the results is to divide the histograms into segments. It is generally agreed that $\rm\Delta BICs$ smaller than two are not significant. We therefore divide the histograms into three segments: $\rm\Delta BIC$ < -2, where the standard dark profile is preferred, -2 < $\rm\Delta BIC$ < 2, where both models are equivalent, and $\rm\Delta BIC$ > 2, where the phenomenological model is preferred. When compared to NFW, the new model is superior in 65% of the cases, the models are equivalent in 15% of the cases, and NFW is better in 20% of the cases (73%-27% for a threshold of zero). When compared to Burkert, the new model is superior in 33% of the cases, the models are equivalent in 37% of the cases, and Burkert is preferred in 30% of the cases (60%-40% for a zero threshold).
@@
Figure 4: Comparing the phenomenological model with established approaches. Left panel: A histogram of $\rm\Delta BIC$ values, where the comparison is relative to NFW. The median value of the histogram is 3.32 $\pm$ 0.75. The error on the median was calculated by Err = $\rm 1.48*MAD/\sqrt{N}$, where N is the number of samples. Right panel: A comparison relative to the Burkert profile. The median value is 0.77 $\pm$ 0.36.
@@
Ultimately, the current analysis suggests that the phenomenon of constant offsets in angular velocity curves is likely a universal property. The hypothesis is supported by the data as it appears comparable to standard dark-halo profiles in terms of model performance. In appendix B, we present a further set of AVCs, spanning a wide range of galaxy sizes, in order to provide the reader with a broader perspective on the nature of AVCs.
@@ 3 Deriving the Total Baryonic Mass
where $\rm\omega_{0}$ and $\rm M$ are the free parameters of the model.
@@
Fig. 5 presents an example of an "outer-AVC" fitting. The model is indicated by a solid red line. The dashed red line, displayed for clarity, represents the inner part of the model, i.e. the region where the Keplerien assumptions are not valid. Ultimately, the fitting procedure determines the values of $\rm\omega_{0}$ and $\rm M$ for each galaxy in the sample.
Having obtained the baryonic masses from the fitting, it would be beneficial to examine their reliability. In Fig. 6 we display the masses by comparing them to those derived from a Burkert fitting. The tight linear correlation demonstrates the viability of the newly-developed application, as it agrees with Burkert across a wide range of masses.
The typical Burkert masses, though, are slightly lower than those derived by the new approach. The median value of the mass difference ($\rm\Delta log(M)$) is $\rm 0.1\pm 0.06$. This is possibly due to the degeneracy between the baryonic disk and the dark halo (Dutton et al., 2005). The Burkert model allows, in some cases, a fit with a negligible baryonic disk, compensated by a maximal dark halo (Li et al., 2020). This could lead to a decrease in the typical Burkert masses relative to the masses obtained by the new method. The opportunity to address the degeneracy with the new application may motivate further exploration.
@@ 4 A Relation between Galaxies and their Host Dark Halos
Figure 6: A scatter plot of the baryonic masses obtained by the model Vs the Burkert-based baryonic masses. A 1:1 dashed-black line is plotted as a reference. The extraction of the Burkert masses is based on publicly available fitting parameters (Li et al., 2020).
@@
Let us take a step further and explicitly present the halo’s mass-density distribution. That is, a halo distribution that induces a gravitational field (and in turn, velocity fields) that are compatible with Eq. 10. The complete derivation is provided in Appendix A. The derivation assumes a spherically symmetric halo, and the derived density profile is given by:
@@
Eq. 11 describes a new dark-halo profile. The halo is characterized by one free parameter, $\rm\omega_{0}$, which represents a constant offset in the data. The free parameter may be replaced with a characteristic density of the form $\rm\rho_{0}=\omega_{0}^{2}/4\pi G$. For a typical value of $\rm\omega_{0}\sim 10^{-16}$ rad/s one would get a typical value of $\rm\rho_{0}\sim 10^{-23}$ kg/m3. This is $\sim 4$ orders of magnitude higher than the average density of dark matter in the universe (Kaplinghat & Turner, 2001; Planck Collaboration et al., 2020), which is expected for a typical value in dense regions.
Most notably, the halo structure is directly connected to the baryonic velocity distributions. This may yield insightful results. For example, if the inner region of the baryonic distribution could be approximated by $\rm v_{bar}\propto r$ then a cored dark halo would appear (i.e. a halo with a constant central density). On the other hand, if the inner region could be associated with a velocity term of the form $\rm v_{bar}\propto\sqrt{r}$ then the inner density profile would be approximated by $\rm\rho_{dark}\propto 1/\sqrt{r}$.
This kind of behaviour may further be explored either by using analytical models or by applying numerical predictions of particular galaxies. In Fig. 7 we employ the second approach to display the new profile’s density distribution of a particular galaxy (NGC-7793). The numerically-derived baryonic velocities (as well as the value of $\rm\omega_{0}$) are the same as in section 2. The profile is plotted alongside the associated NFW and Burkert profiles. Their parameters were taken from Li et al. (2020) to adjust these profiles to the current galaxy. Interestingly, the inner slope of the new density profile lies between Burkert and NFW with a value of $\rm-0.7$. The surrounding regions roughly appear to resemble the established profiles.
The coupling between dark and baryonic distributions in spiral galaxies has been extensively discussed in the literature (Persic et al., 1996; Salucci & Burkert, 2000; Swaters et al., 2009; McGaugh, 2014). The contribution of the current study is the introduction of an explicit relation that connects these two distributions.

### v2

@@ Abstract
The well-known discrepancy in galactic rotation curves refers to the mismatch between observed rotational velocities and the velocities predicted by baryonic matter. In this study we investigate a potential pattern in the discrepancy, which may point to an underlying pattern in dark-halo distributions. By looking at rotational-velocity curves from an alternate perspective, the angular-velocity curves, it appears that the observed angular velocities and their corresponding baryonic predictions differ by a constant shift. That is, the discrepancy may be reduced to a constant angular-velocity term, independent of the radius. We test the generality of the suggested property by analysing 143 high-quality rotation curves. The property appears significant as it performs equally well (or better) than well-established models. Compared to a Burkert dark-halo profile, it is preferred in 60% of the cases, relative to a Navarro–Frenk–White profile (NFW), it is superior in 73% of the cases, and relative to Modified Newtonian Dynamics (MOND), it exhibits similar performance, being favoured in 50% of the cases. Next, by including the new phenomenological property within the dynamical equations, we find an explicit expression for the dark-halo profile. The new single-parameter profile is characterised by an interesting property: it is intrinsically related to the baryonic distribution. Thus, information regarding the cuspy or cored nature of a particular dark halo, according to this profile, is encoded (and explicitly determined) by the respective baryonic behaviour.
@@
pubyear: 2025pagerange: On the Presence of Angular-Velocity Offsets in Disk GalaxiesE
@@ 1 Introduction
One of the most useful tools to analyse the dynamics of disk galaxies and derive their mass distributions is a galactic rotation curve (Bosma, 1981; Sofue and Rubin, 2001). A rotation curve (RC) presents the circular velocities in a galaxy as a function of the distance from the galaxy’s center. Historically, the discrepancy revealed in RCs is known as the first robust evidence for dark matter (Babcock, 1939; Salpeter, 1978; Rubin et al., 1978, 1980; Carignan, 1985; Sancisi, 1987). Illustrations of the discrepancy (i.e. the mismatch between baryonic predictions and observations) are shown in the upper panels of Fig. 1.
@@
Although the baryonic distributions are less dominant than their accompanying dark halos (Kuhlen et al., 2012; Reddick et al., 2013), their impact on the overall dynamics seems to be important. A notable result of such a kind was found by Sancisi (2004), who showed that the features in baryonic distributions closely correlate with features in rotation curves. Since the velocities in rotation curves reflect the total gravitational field (i.e., of the overall mass distribution), such a correlation highlights the important role of baryons in describing the total dynamics.
@@
The objective of the current paper is to demonstrate, establish, and explore the implications of the constant-offsets property. A natural outcome of the property is an explicit relation that connects dark and baryonic distributions in disk galaxies. This can be achieved by phenomenologically integrating the offsets into the dynamics. The resulting relation can be interpreted as a new data-driven dark-halo profile.
@@
The next sections are organized as follows: Section 2 presents the new concept of angular-velocity curves and explores the universality of constant offsets. Section 3 calculates the baryonic masses of 128 disk galaxies using their observed rotation curves, and without additional assumptions. Section 4 derives a relation that connects baryonic distributions with their host dark halos, or equivalently, introduces a new dark-halo profile. In Section 5, we discuss potential interpretations of the observed property, conclude, and propose future directions.
@@ 2 Angular-Velocity Curves
[…]
@@
Our aim is to compare the new phenomenological relation (Eq. 1) with established approaches. Standard dark-halo profiles and the Modified Newtonian Dynamics model (MOND, Milgrom, 1983) already produce reliable fits to the data, so they may act as a baseline. If the new model (baryons + offset) could produce comparable results to these baseline models, it may be considered a viable approach. This would support the universality of angular-velocity offsets in disk galaxies.
@@
Figure 3: A visual representation of the sample’s best-fit model parameters. Left panel: A scatter plot of the two parameters. There is a weak but not significant anti-correlation between them ($\rm\rho_{pearson}=-0.1$, $p_{\rm value}=0.26$). Middle panel: Histogram of M/L values. The median value of log(M/L) is $\rm 0.06\pm 0.04$. There are four outliers in the data with $\rm M/L<0.1$, but within uncertainties, they are also consistent with higher values. Right Panel: Histogram of $\rm\omega_{0}$ values. The median value of $\rm log(\omega_{0})$ is $15.97\pm 0.51$. Eight points with $\rm\omega_{0}<0$ are not shown, due to the logarithmic nature of the plot, but within uncertainties, they are also consistent with positive values.
Dark-halo models usually include three free parameters: $\rm M/L$ for the baryonic component and two more for the dark halo. In the current analysis we focus on two dark halo profiles: the cosmology-driven Navarro–Frenk–White (NFW, Navarro et al., 1997) and the data-driven Burkert (Burkert, 1995). In order to quantify the goodness of fit of the dark-halo models we rely on the catalog of Li et al. (2020). For each galactic rotation curve, the catalog specifies the best-fit parameters and the corresponding $\chi^{2}$.
MOND includes one free parameter per galaxy: the $\rm M/L$ of the baryonic component. The critical acceleration $\rm a_{0}$, which is a universal constant in MOND theory, is treated as a global parameter. In order to quantify MOND’s goodness of fit we also perform an MCMC simulation. The technical details are provided in Appendix A.
Since we prefer to use exactly the same data-sets when testing the competing models (i.e., NFW, Burkert, MOND and the current one), we convert the angular velocities of Eq.  1 into actual velocities, as follows:
@@
where $\rm M/L$ is the stellar-disk mass-to-light ratio, $v_{\rm disk}(r)$ is the normalized stellar-disk contribution and $v_{\rm gas}(r)$ is the gaseous-disk contribution. The normalized stellar-disk and the gaseous-disk contributions are included in the SPARC database as data files. They were numerically derived from the corresponding light and neutral hydrogen distributions. More information on the process can be found in Sanders and McGaugh (2002).
Baryonic contributions are in fact included in all galactic rotation-curve models. In the general case, they may also contain a bulge term. However, in the current study, to avoid the complexity of adding an extra $\rm M/L$ parameter to the fittings (and the statistical analysis), or alternatively to avoid extra sensitivity to stellar population synthesis models, only galaxies with negligible bulge contributions were selected. This leaves us with 143 galaxies.
The next step is to apply priors to the parameters. For the $\rm M/L$ parameter we impose a prior of the form $\rm M/L$ > 0 to ensure that it maintains its physical meaning. For $\rm\omega_{0}$ we apply a non-informative uniform prior since no earlier knowledge is available. We then run the MCMC simulation through $10,000$ steps, using a Metropolis-Hastings algorithm and a burn-time of $800$ steps. To ensure the simulation’s convergence, we apply a Gelman-Rubin test and make sure that the potential scale reduction factors are smaller than $1.2$ (Hogg and Foreman-Mackey, 2018; Gelman and Rubin, 1992). The MCMC outputs are the estimated parameters and the corresponding $\rm\chi^{2}$ of each fit. An example of an MCMC fitting (converted to angular velocities for convenience) is given in Fig. 2.
@@
Another aspect would be the typical values of the parameters. Fig. 3 displays the distributions of the parameters. Table 1 provides the values, their errors, and additional statistical information. The median value of $\rm\omega_{0}$ is $\sim 10^{-16}$ [rad/s], with a range of almost two orders of magnitude. The median value of $\rm M/L$ is $\sim\rm 1$ $\rm M_{\odot}/L_{\odot}$], which is somewhat higher than expectations (McGaugh and Schombert, 2014). Appendix B tackles this concern by applying an additional fitting procedure, using different priors. Finally, there is a weak but not significant anti-correlation between the parameters ($\rm\rho_{pearson}=-0.1$, $p_{\rm value}=0.26$).
Table 1: Best-fit parameters and fitting information of Section 2
Galaxy $\rm M/L$ $\rm\omega_{0}$ $\rm n$ $\rm\chi^{2}$ $\rm p_{value}$
$\rm M_{\odot}/L_{\odot}$] $\rm[10^{-16}\,rad/s]$
@@
With both the MCMC and the catalog results in hand, the relative performance of the different models could be quantified. That is, the performance of the phenomenological model with respect to the established models. For this purpose we employ the Bayesian Information Criterion (BIC). BIC is a criterion for model selection that accounts for the trade-off between a model’s fit and its complexity. It is suitable for comparing non-nested models when the data are assumed to be independent and normally distributed, with the noise being well characterized by measurement errors. The BIC score of a given model (with respect to a particular dataset) can be written as:
@@
We calculate the BIC difference (i.e. the $\rm\Delta BIC$ between NFW/Burkert/MOND and the phenomenological model) for the entire sample. A positive value indicates that the phenomenological model performs better than the baseline model and vice versa.
@@
On the central panel of Fig. 4, we plot a similar histogram, with the comparison being relative to Burkert’s profile. In this case, the $\rm\Delta BIC$ test suggests that the models are equally credible, with slightly better results for the phenomenological model (median value of 0.77 $\pm$ 0.36). Flawed results were excluded as well (14 galaxies where Burkert fails, 20 where the phenomenological model fails, and 16 where both models fail).
@@
Figure 4: Comparing the phenomenological model with established approaches. Left panel: A histogram of $\rm\Delta BIC$ values, where the comparison is relative to NFW. The median value of the histogram is 3.32 $\pm$ 0.75. The error on the median was calculated by Err = $\rm 1.48*MAD/\sqrt{N}$, where N is the number of samples. Central panel: A comparison relative to the Burkert profile. The median value is 0.77 $\pm$ 0.36. Right panel: A comparison relative to MOND. The median value is -0.19 $\pm$ 0.68.
In the right panel of Fig. 4 we compare the current model to MOND. The $\rm\Delta BIC$ test suggests that the present model and MOND are equally credible, with a median value of -0.19 $\pm$ 0.68. It is necessary to note that there are several possible versions of MOND (e.g., Li, Pengfei et al. (2018); Wang and Chen (2021)). Appendix A examines a few versions and their respective results. The adopted (main-text) result is based on Li, Pengfei et al. (2018). Flawed results were excluded as well (22 galaxies where MOND fails, 15 where the phenomenological model fails, and 31 where both models fail).
While the number of excluded galaxies appears to be high, the essential aspect here is that the numbers are comparable. In some cases the established models fail, whereas in others it is the phenomenological model that has difficulty fitting the data. The challenges in fitting the data depend on the accuracy of the data itself but also on the modeling of the baryonic component. These difficulties may include the extraction of surface photometry (Lelli et al., 2016), the bulge-disk decomposition (Simard et al., 2011) and the irregular morphologies of low surface-brightness galaxies.
Another way to look at the results is to divide the histograms into segments. It is generally agreed that $\rm\Delta BICs$ smaller than two are not significant. We therefore divide the histograms into three segments: $\rm\Delta BIC$ < -2, where the established model is preferred, -2 < $\rm\Delta BIC$ < 2, where both models are equivalent, and $\rm\Delta BIC$ > 2, where the phenomenological model is preferred.
When compared to NFW, the new model is superior in 65% of the cases, the models are equivalent in 15% of the cases, and NFW is better in 20% of the cases (73%-27% for a threshold of zero). When compared to Burkert, the new model is superior in 33% of the cases, the models are equivalent in 37% of the cases, and Burkert is preferred in 30% of the cases (60%-40% for a zero threshold). Finally, when compared to MOND, the new model is superior in 36% of the cases, the models are equivalent in 25% of the cases, and MOND is preferred in 39% of the cases (50%-50% for a zero threshold).
@@
Ultimately, the current analysis suggests that the phenomenon of constant offsets in angular velocity curves is likely a universal property. The hypothesis is supported by the data as it appears comparable to standard dark-halo profiles and MOND. In Appendix E, we present a further set of AVCs, spanning a wide range of galaxy sizes, in order to provide the reader with a broader perspective on the nature of AVCs.
@@ 3 Deriving the Total Baryonic Mass
where $\rm\omega_{0}$ and $\rm M$ are the free parameters of the model. Note that this model does not rely on measured quantities (unlike Eq. 3 for example) and is therefore insensitive to the difficulties associated with their calculation.
@@
Fig. 5 presents an example of an "outer-AVC" fitting. The model is indicated by a solid red line. The dashed red line, displayed for clarity, represents the inner part of the model, i.e. the region where the Keplerien assumptions are not valid. Ultimately, the fitting procedure determines the values of $\rm\omega_{0}$ and $\rm M$ for each galaxy in the sample. The values of $\omega_{0}$ are not precisely identical to those obtained in Section 2 since the fitting approach is different in this case.
Finally, the derived baryonic masses are given in Table 2. The median value of the population is $\rm 1.66\pm 1.1\ [10^{10}\ M_{\odot}]$, with the masses spanning $\rm 4$ orders of magnitude. The full table is available online (see Data Availability).
@@ 4 A Relation between Galaxies and their Host Dark Halos
Table 2: Best-fit parameters and fitting information of Section 3.
Galaxy $\rm M$ $\rm\omega_{0}$ $\rm n$ $\rm\chi^{2}$ $\rm p_{value}$
$\rm 10^{10}\ M_{\odot}$] $\rm[10^{-16}\,rad/s]$
NGC 3949 $1.23^{+0.59}_{-0.46}$ $3.77^{+1.14}_{-1.20}$ 3 0.0087 0.92
NGC 4100 $7.59^{+0.33}_{-0.38}$ $0.63^{+0.08}_{-0.06}$ 13 6.64 0.83
NGC 4214 $0.34^{+0.14}_{-0.12}$ $1.58^{+0.93}_{-0.91}$ 7 0.55 0.99
Note: $\rm M$ denotes the total baryonic mass, $\rm\omega_{0}$ is the angular-velocity offset, $n$ is the number of data points in the outer part of the rotation curve, $\chi^{2}$ is the best-fit sum of normalized residuals, and $\rm p_{value}$ represents the corresponding probability. The full table is available online (see Data Availability Section).
@@
Let us take a step further and explicitly present the halo’s mass-density distribution. That is, a halo distribution that induces a gravitational field (and in turn, velocity fields) that are compatible with Eq. 10. The complete derivation is provided in Appendix C. The derivation assumes a spherically symmetric halo, and the derived density profile is given by:
@@
Eq. 11 describes a new dark-halo profile. The halo is characterized by one free parameter, $\rm\omega_{0}$, which represents a constant offset in the data. The free parameter may be replaced with a characteristic density of the form $\rm\rho_{0}=\omega_{0}^{2}/4\pi G$. For a typical value of $\rm\omega_{0}\sim 10^{-16}$ rad/s one would get a typical value of $\rm\rho_{0}\sim 10^{-23}$ kg/m3. This is $\sim 4$ orders of magnitude higher than the average density of dark matter in the universe (Kaplinghat and Turner, 2001; Planck Collaboration et al., 2020), which is expected for a typical value in dense regions.
Furthermore, the halo structure is directly connected to the baryonic velocity distributions. This may yield insightful results. For example, if the inner region of the baryonic distribution could be approximated by $\rm v_{bar}\propto r$ then a cored dark halo would appear (i.e. a halo with a constant central density). On the other hand, if the inner region could be associated with a velocity term of the form $\rm v_{bar}\propto\sqrt{r}$ then the inner density profile would be approximated by $\rm\rho_{dark}\propto 1/\sqrt{r}$.
This kind of behaviour may further be explored either by using analytical models or by applying numerical models of particular galaxies. In Fig. 6 we employ the second approach and display the new profile’s density distribution for a particular galaxy (NGC-7793). The value of $\rm\omega_{0}$ (and the numerically-derived baryonic velocities) are the same as in Section 2. The profile is plotted alongside the associated NFW and Burkert profiles, with their parameters taken from Li et al. (2020). Interestingly, the inner slope of the new density profile lies between Burkert and NFW with a value of $\rm-0.7$. The surrounding regions roughly appear to resemble the established profiles.
The coupling between dark and baryonic distributions in spiral galaxies has been extensively discussed in the literature (Persic et al., 1996; Salucci and Burkert, 2000; Swaters et al., 2009; McGaugh, 2014). The current study introduces an explicit relation between these two distributions.